The demand function for a dress is y = 47−3p and the supply function is y = 6+1p, where p is the price of a dress and y is the quantity of dresses. The government introduces two policies: first, it fixes the price of a dress at 7 per unit; second, it promises to pay producers a subsidy for each dress produced, so as to make supply equal demand. How much would the subsidy per dress have to be?

Select one:

a. 10.2

b. -13

c. None of the answers is correct.

d. 13

e. 0

f. 7
Answer #1

Correct Answer:

C

Working note:

At equilibrium:

47-3P = 6+P

41 = 4P

P = 41/4 = 10.25

Price fixed = 7

So, subsidy given = 10.25 - 7 = 3.25

So, subsidy to be given is 3.25 so that equilibrium is maintained.